# Scattering wave packets of hadrons in gauge theories: Preparation on a quantum computer

### v2

@@ Acknowledgment
We thank Brian Neyenhuis and Jenni Strabley from Quantinuum for their generosity and support, and Microsoft Azure and the Division of Information Technology at the University of Maryland (UMD IT) for providing the infrastructures to access the quantum emulator and hardware via Cloud Services. We are grateful to Niklas Mueller for his early effort in ensuring access to the Azure Cloud subscription via UMD IT, and for the help in acquisition of some of the quantum-computing credits used in this project. We further thank Connor Powers for assistance with aspects of running our codes on a local computing machine and using Azure Cloud Services. We further acknowledge useful discussions with Navya Gupta, Jesse Stryker, and Francesco Turro. The numerical calculations presented in this work are made possible by the extensive use of Qiskit Qiskit contributors (2023) and Jupyter Notebook Kluyver et al. (2016) software applications within the Conda Anaconda Software Distribution environment. Z.D., C-C.H., and S.K. acknowledge funding by the U.S. Department of Energy (DOE), Office of Science, Early Career Award (award no. DESC0020271), and the DOE, Office of Science, Office of Nuclear Physics, via the program on Quantum Horizons: QIS Research and Innovation for Nuclear Science (award no. DE-SC0021143). Z.D. and C-C.H. are grateful for the hospitality of Perimeter Institute where part of this work was carried out. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science, and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities. This research was also supported in part by the Simons Foundation through the Simons Foundation Emmy Noether Fellows Program at Perimeter Institute. S.K. further acknowledges support by the U.S. DOE, Office of Science, Office of Nuclear Physics, InQubator for Quantum Simulation (IQuS) (award no. DE-SC0020970), and by the DOE QuantISED program through the theory consortium “Intersections of QIS and Theoretical Particle Physics” at Fermilab (Fermilab subcontract no. 666484). This work was also supported, in part, through the Department of Physics and the College of Arts and Sciences at the University of Washington.
Appendix A Numerical verification of the ansatz for 10-site theories
@@
Here, we provide the ansatz-optimization results for the 10-site theory. For a lattice with 10 staggered sites, there are 5 discrete momentum sectors in the Brillouin zone, $k\in\widetilde{\Gamma}=\{-\frac{2\pi}{5},-\frac{\pi}{5},0,\frac{\pi}{5},\frac{2\pi}{5}\}$. Figure 11 shows the fidelity between the optimized |$\ket{k}$ and the exact-diagonalized momentum eigenstates for $k=0,\frac{\pi}{5},\frac{2\pi}{5}$ in both the $Z_{2}$ and the $U(1)$ LGTs, with a cutoff of $\Lambda=1$ in the $U(1)$ case. Results for negative $k$ values are identical to those for the corresponding positive values and are not shown.
@@
Similar to Fig. 5 for the 6-site theory, the region where the low-lying purely non-mesonic states are present is marked to the left side of the cyan line. The energy required to excite such states from the interacting ground state is proportional to the system size. Compared to the 6-site theory, such a region is reduced as in the 10-site theory as expected.
In the 10-site theory, both |$\ket{k=0}$ and |$\ket{k=\pm\frac{\pi}{5}}$ can be optimized with $1-F=\mathcal{O}(10^{-3})$ for the majority of the parameter space. For parameters tuned toward the weak-coupling regime (i.e., $\epsilon,m_{f}\to 0$ for a finite $m_{f}/\epsilon$), the performance is seen to be worse. For $k=\pm\frac{2\pi}{5}$, the fidelity between the optimized state and the exact-diagonalized state is around 0.90 for most of the parameter space in both theories. It is generally observed that the $k$-momentum optimization becomes less accurate with $k$ closer to the boundary of the Brillouin zone, $\tilde{\Gamma}$. Therefore, if the wave packet is centered further away from the edge of the Brillouin zone, the Gaussian profile of the wave packet will suppress the effect of these lower-fidelity high-momentum states in the full wave-packet state.
@@
Figure 11: The $k$-optimization heat map for the 10-site theories. Panels (a) and (b) refer to the $Z_{2}$ LGT and the $U(1)$ LGT (with a cutoff of $1$), respectively. The region left to the cyan contour is where the purely non-mesonic states are less energetic than the single-meson |$\ket{k}$ state. The optimization works well for |$\ket{k=0}$ and |$\ket{k=\pm\frac{\pi}{5}}$, but is less effective for |$\ket{k=\pm\frac{2\pi}{5}}$.
@@ Appendix D Wave packets with quantum-hardware emulators
In this appendix, we investigate the effect of statistical shot noise compared with the hardware-specific sources of error. To this aim, we present the same quantities as in Fig. 10 of the main text in Fig. 15, removing the exact and ideal-circuit results, while adding Qiskit’s AER simulator Qiskit contributors (2023) with 500 measurement shots (as used to obtain the hardware results). The AER simulator provides values affected by only the statistical shot noise. It is observed that the effect of shot noise is generally small, and the hardware-specific error likely contributes more to the disagreement between the statevector-evolution results and the hardware results.
It is also valuable to compare the values obtained by the Quantinuum H1-1 emulator, which incorporates both the statistical shot noise and hardware-specific noise. The latter is modeled to faithfully emulate the hardware performance. These results are also included in the plots in Fig. 15. There is a reasonable agreement between the emulator and the hardware results, indicating that the noisy emulator describes the quantum hardware faithfully. The uncertainties on the AER simulator and the Quantinuum H1-1 emulator results are obtained using the same bootstrap procedure described for the hardware results in Sec. III.2.2. Since we have limited hardware access but less restriction on emulator access, we can perform further studies with this emulator. As an example, to verify the behavior of the wave-packet preparation circuit as a function of other parameters of the wave packet beside its width, we have obtained the local fermionic density $\chi_{n}$, defined in Eq. (26), for different values of $\mu$ and $k_{0}$. The results are plotted in Fig. 16. 500 shots are used for each parameter set $(\sigma,\mu,k_{0})$, and the physical error mitigation mentioned in Sec. III.2.2 is applied to the Quantinuum H1-1 emulator results. The results are in agreement with the expectations: changing $\mu$ corresponds to changing the center of the wave packet in position space, while changing $k_{0}$ corresponds to changing the wavenumber in position space.
@@ Appendix E Physical Hilbert space of the 6-site theories
In this appendix, we list all 40 (38) physical configurations that satisfy the Gauss’ law [Eq. (2)] in the 6-site $Z_{2}$ LGT ($U(1)$ LGT) in 1+1 D with one flavor of staggered fermions. The assigned labels of these configurations (i.e., physical basis states) are used as the x-axis in Figs. 6, 9, 10, and 15. As mentioned in the main text, the physical Hilbert space is restricted to the subspace with $Q=N/2$, where the fermion-number operator $Q$ is defined in Sec. II.1. For the $U(1)$ LGT, the electric-field basis is cut off at the value $\Lambda=1$, allowing only three possible electric-field values ($0,\pm 1$) at each link.
Physical-configuration label |$\ket{0}_{f}$ |$\ket{0}_{b}$ |$\ket{1}_{f}$ |$\ket{1}_{b}$ |$\ket{2}_{f}$ |$\ket{2}_{b}$ |$\ket{3}_{f}$ |$\ket{3}_{b}$ |$\ket{4}_{f}$ |$\ket{4}_{b}$ |$\ket{5}_{f}$ |$\ket{5}_{b}$
@@
Physical-configuration label |$\ket{0}_{f}$ |$\ket{0}_{b}$ |$\ket{1}_{f}$ |$\ket{1}_{b}$ |$\ket{2}_{f}$ |$\ket{2}_{b}$ |$\ket{3}_{f}$ |$\ket{3}_{b}$ |$\ket{4}_{f}$ |$\ket{4}_{b}$ |$\ket{5}_{f}$ |$\ket{5}_{b}$

### v3

@@ Acknowledgment
We thank Brian Neyenhuis and Jenni Strabley from Quantinuum for their generosity and support, and Microsoft Azure and the Division of Information Technology at the University of Maryland (UMD IT) for providing the infrastructures to access the quantum emulator and hardware via Cloud Services. We are grateful to Niklas Mueller for his early effort in ensuring access to the Azure Cloud subscription via UMD IT, and for the help in acquisition of some of the quantum-computing credits used in this project. We further thank Connor Powers for assistance with aspects of running our codes on a local computing machine and using Azure Cloud Services. We further acknowledge useful discussions with Navya Gupta, Jesse Stryker, and Francesco Turro. The numerical calculations presented in this work are made possible by the extensive use of Qiskit [146] and Jupyter Notebook [147] software applications within the Conda [148] environment. Z.D., C-C.H., and S.K. acknowledge funding by the U.S. Department of Energy (DOE), Office of Science, Early Career Award (award no. DESC0020271), and the DOE, Office of Science, Office of Nuclear Physics, via the program on Quantum Horizons: QIS Research and Innovation for Nuclear Science (award no. DE-SC0021143). Z.D. and C-C.H. are grateful for the hospitality of Perimeter Institute where part of this work was carried out. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science, and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities. This research was also supported in part by the Simons Foundation through the Simons Foundation Emmy Noether Fellows Program at Perimeter Institute. S.K. further acknowledges support by the U.S. DOE, Office of Science, Office of Nuclear Physics, InQubator for Quantum Simulation (IQuS) (award no. DE-SC0020970), and by the DOE QuantISED program through the theory consortium “Intersections of QIS and Theoretical Particle Physics” at Fermilab (Fermilab subcontract no. 666484). This work was also supported, in part, through the Department of Physics and the College of Arts and Sciences at the University of Washington.
Appendix A Numerical verification of the ansatz for $U(1)$ LGT
@@
Supplementing the results of Sec. 2.3, in this Appendix we provide the ansatz-optimization results for the $U(1)$ LGT in 1+1 D. We consider the 6-site and the 10-site theories using a cutoff of $\Lambda=1$. Figure 11 shows the fidelity between the optimized |$\ket{k}$ and the exact-diagonalized momentum eigenstates for the same Brillouin zone as in Fig. 5. Results for negative $k$ values are identical to those for the respective positive values and are not shown.
@@
Similar to Fig. 5 for the $Z_{2}$ LGT, the region where the low-lying purely non-mesonic states are present is marked to the left side of the cyan line. The energy required to excite such states from the interacting ground state is proportional to the system size. We observed the same reduction of the non-mesonic excitation regime from 6 sites to 10 sites as in the case of $Z_{2}$ LGT. Finally, we plot in Fig. 12 the probability distribution of each of the physical configurations in the $U(1)$ theory with parameters $m_{f}=1$ and $\epsilon=1.0$. Each $k$-momentum state created by the ansatz recovers the exact-diagonalized state probabilities.
@@
Figure 11: The $k$-optimization heat map for the $U(1)$ LGT. Panels (a) and (b) refer to the 6-site and the 10-site theories, respectively (with a cutoff of $\Lambda=1$). The region left to the cyan contour is where the purely non-mesonic states are less energetic than the single-meson |$\ket{k}$ state. The optimization works well for the 6-site theory and |$\ket{k=0}$ and |$\ket{k=\pm\frac{\pi}{5}}$ in the 10-site theory, but is less effective for |$\ket{k=\pm\frac{2\pi}{5}}$, where fidelities of only $90\%$ are reached for most of the parameter space.
Figure 12: Optimization results for $k=0\leavevmode\nobreak\ (\frac{\pi}{3})$ momentum eigenstates are shown in the left (right) panel for the case of 6-site $U(1)$ LGT with $m_{f}=1$, $\epsilon=1.0$, and cutoff $\Lambda=1$. The physical basis-state probabilities the of $k=-\frac{\pi}{3}$ state are identical to $k=\frac{\pi}{3}$ and are not shown. The physical basis states of the 6-site $U(1)$ theory are listed in Tables 3.
@@ Appendix B Efficient-circuit parameters
The analysis above considers the theory with $m=1.0$, $\epsilon=-0.3$. Additionally, the effect of $\theta_{c}$ truncation for the same wave-packet parameters but different $m_{f}$ and $\epsilon$ values is shown in Fig. 15. Qualitatively, there are more terms surviving the truncation in the low-mass regime.
Figure 15: The number of mesonic creation operators using the wave-packet parameters ($\sigma=\frac{\pi}{6},\mu=3,k_{0}=0$) but with a wide range of $m_{f}$, $\epsilon$ values is plotted. The left (right) panels corresponds to $\theta_{c}=0.1(0.05)$. The maximum number of terms is $N^{2}$ (36 for the 6-site theory). It is observed that more terms survive the $\theta_{c}$ truncation in the low-mass regime.
@@ Appendix D Wave packets with quantum-hardware emulators
In this appendix, we investigate the effect of statistical shot noise compared with the hardware-specific sources of error. To this aim, we present the same quantities as in Fig. 10 of the main text in Fig. 17, removing the exact and ideal-circuit results, while adding Qiskit’s AER simulator [146] with 500 measurement shots (as used to obtain the hardware results). The AER simulator provides values affected by only the statistical shot noise. It is observed that the effect of shot noise is generally small, and the hardware-specific error likely contributes more to the disagreement between the statevector-evolution results and the hardware results.
It is also valuable to compare the values obtained by the Quantinuum H1-1 emulator, which incorporates both the statistical shot noise and hardware-specific noise. The latter is modeled to faithfully emulate the hardware performance. These results are also included in the plots in Fig. 17. There is a reasonable agreement between the emulator and the hardware results, indicating that the noisy emulator describes the quantum hardware faithfully. The uncertainties on the AER simulator and the Quantinuum H1-1 emulator results are obtained using the same bootstrap procedure described for the hardware results in Sec. 3.2.2. Since we have limited hardware access but less restriction on emulator access, we can perform further studies with this emulator. As an example, to verify the behavior of the wave-packet preparation circuit as a function of other parameters of the wave packet beside its width, we have obtained the local fermionic density $\chi_{n}$, defined in Eq. (26), for different values of $\mu$ and $k_{0}$. The results are plotted in Fig. 18. 500 shots are used for each parameter set $(\sigma,\mu,k_{0})$, and the physical error mitigation mentioned in Sec. 3.2.2 is applied to the Quantinuum H1-1 emulator results. The results are in agreement with the expectations: changing $\mu$ corresponds to changing the center of the wave packet in position space, while changing $k_{0}$ corresponds to changing the wavenumber in position space.
@@ Appendix E Physical Hilbert space of the 6-site theories
In this appendix, we list all 40 (38) physical configurations that satisfy the Gauss’ law [Eq. (2)] in the 6-site $Z_{2}$ LGT ($U(1)$ LGT) in 1+1 D with one flavor of staggered fermions. The assigned labels of these configurations (i.e., physical basis states) are used as the x-axis in Figs. 6, 9, 10, and 17. As mentioned in the main text, the physical Hilbert space is restricted to the subspace with $Q=N/2$, where the fermion-number operator $Q$ is defined in Sec. 2.1. For the $U(1)$ LGT, the electric-field basis is cut off at the value $\Lambda=1$, allowing only three possible electric-field values ($0,\pm 1$) at each link.
Physical- |$\ket{0}_{f}$ |$\ket{0}_{b}$ |$\ket{1}_{f}$ |$\ket{1}_{b}$ |$\ket{2}_{f}$ |$\ket{2}_{b}$ |$\ket{3}_{f}$ |$\ket{3}_{b}$ |$\ket{4}_{f}$ |$\ket{4}_{b}$ |$\ket{5}_{f}$ |$\ket{5}_{b}$
configura-
tion label
@@
Physical- |$\ket{0}_{f}$ |$\ket{0}_{b}$ |$\ket{1}_{f}$ |$\ket{1}_{b}$ |$\ket{2}_{f}$ |$\ket{2}_{b}$ |$\ket{3}_{f}$ |$\ket{3}_{b}$ |$\ket{4}_{f}$ |$\ket{4}_{b}$ |$\ket{5}_{f}$ |$\ket{5}_{b}$
configura-
tion label